OpenStudy (anonymous):

Verify identity: sin4t=4sintcost(1-2sinˆ2t) Got as far as: LHS sin(2x2t)= 2sin2tcos2t= and i'm stuck....

OpenStudy (mertsj):

Very hard to tell what the original problem is.

OpenStudy (mertsj):

\[\sin 4t=4\sin t \cos t(1-2\sin ^2t)\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is the original problem

OpenStudy (mertsj):

\[\sin 4t=\sin (2t+2t)=\sin 2t \cos 2t+\cos 2t \sin 2t=2\sin 2t \cos 2t=\] \[2(2\sin t \cos t)(1-2\sin ^2t)=4\sin t \cos t(1-2\sin ^2t)\]